OpenStudy (anonymous):

Write as the sum of the product cos11x+cos2x

OpenStudy (anonymous):

\[\cos(x)+\cos(y)=2\cos(\frac{x+y}{2})\cos(\frac{x-y}{2})\] if i remember correctly b